# Pólya’s conjecture for thin products

### v2

@@ 2. Some preparations
2.1. Two term inequalities for the eigenvalue counting functions and the Riesz means
@@
where $|\partial\Omega|$ is the surface area of $\partial\Omega$. Although Weyl’s conjecture was not proven in its full generality, R. Seeley ([31], [32]) proved a weaker version, namely both eigenvalue counting functions satisfy
$\mathcal{N}^{D/N}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}+\mathrm{O}(\lambda^{\frac{d-1}{2}}),\qquad\text{as\ }\lambda\to\infty,$
@@
We also need two-term inequalities for the Riesz mean that sharpen Laptev’s inequalities (1.6) and (1.8). For the Dirichlet case, R. L. Frank and S. Larson ([10]) proved that for any bounded domain $\Omega$ in $\mathbb{R}^{d}$ ($d\geq 2$) with Lipschitz boundary and any $\gamma\geq 1$,
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}})$
@@
(2.8) $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}\leq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{5}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}.$
For the Neumann case, R. L. Frank and L. Geisinger ([9]) proved that for any bounded domain $\Omega$ in $\mathbb{R}^{d}$ ($d\geq 2$) with $C^{1}$ boundary and any $\gamma\geq 1$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}+\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}})$
as $\lambda\to\infty$. As a consequence, for fixed $\gamma$, there exists a positive constant $C_{2}(\Omega)$ such that if $\lambda>C_{2}(\Omega)$, one has
@@
2.2. Strict inequalities for Riesz means
Another ingredient in our proof is the inequality (1.6) of A. Laptev, which follows from
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
By carefully analyzing his proof, one can see that the inequality is strict. We include the proof for reader’s convenience.
Proposition 2.1.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$, then for any $\lambda>0$, one has
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))<L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Proof.
We first sketch Laptev’s proof ([21, Theorem 2.1]). Let $\{\varphi_{k}\}_{k=1}^{\infty}$ be the $L^{2}$ normalized eigenfunctions associated to $\{\lambda_{k}(\Omega)\}_{k=1}^{\infty}$. Then
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq(2\pi)^{-d}\int_{\mathbb{R}^{d}}(\lambda-|\xi|^{2})_{+}\cdot\sum_{\lambda_{k}(\Omega)<\lambda}|\hat{\varphi}_{k}(\xi)|^{2}\mathrm{d}\xi.$
Let
$e_{\xi}(x)=\begin{cases}e^{-ix\cdot\xi},&\qquad x\in\Omega,\\ 0,&\qquad x\notin\Omega.\end{cases}$
Then
$\sum_{\lambda_{k}(\Omega)<\lambda}|\hat{\varphi}_{k}(\xi)|^{2}=\sum_{\lambda_{k}(\Omega)<\lambda}\langle e_{\xi},\varphi_{k}\rangle^{2}_{L^{2}(\Omega)}\\ \leq\|e_{\xi}\|^{2}_{L^{2}(\Omega)}=|\Omega|.$
So
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq(2\pi)^{-d}|\Omega|\int_{\mathbb{R}^{d}}(\lambda-|\xi|^{2})_{+}\mathrm{d}\xi=L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Now assume by contradiction that $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))=L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}$ for some $\lambda$, then
$\sum_{\lambda_{k}(\Omega)<\lambda}\langle e_{\xi},\varphi_{k}\rangle^{2}_{L^{2}(\Omega)}=\|e_{\xi}\|^{2}_{L^{2}(\Omega)},\qquad\forall|\xi|^{2}<\lambda.$
Take $N$ such that $\lambda_{N}(\Omega)<\lambda\leq\lambda_{N+1}(\Omega)$, then one has $e_{\xi}\in\mathrm{span}\{\varphi_{1},\cdots,\varphi_{N}\}$ for any $\xi$ with $|\xi|^{2}<\lambda$. So if we take $N+1$ different points $\{\xi_{1},\cdots,\xi_{N+1}\}$ with $|\xi_{i}|^{2}<\lambda$, $\forall 1\leq i\leq N+1$, then there exists $(c_{1},\cdots,c_{N+1})\neq(0,\cdots,0)$ such that
$\sum_{i=1}^{N+1}c_{i}e_{\xi_{i}}(x)=0,\qquad\forall x\in\Omega.$
By the analyticity of $e^{-ix\cdot\xi}$, one has
$\sum_{i=1}^{N+1}c_{i}e^{-ix\cdot\xi_{i}}=0,\qquad\forall x\in\mathbb{R}^{d},$
which is a contradiction. ∎
Together with the fact
(2.10) $\int^{\infty}_{0}(z-\lambda-t)_{+}t^{\gamma-2}\mathrm{d}t=\frac{(z-\lambda)^{\gamma}_{+}}{\gamma(\gamma-1)},\qquad\forall\gamma>1,$
we immediately get
Corollary 2.2.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$, then for any $\lambda>0$ and $\gamma\geq 1$, one has
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}<L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}.$
In particular, for any $0<a<b$, one has
(2.11) $K(a,b,\Omega):=\inf_{a\leq\lambda\leq b}\frac{L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}}{\lambda^{\gamma+\frac{d-1}{2}}}>0.$
We also need the Neumann version of the same (strict) inequality, which is essentially proved by A. Laptev (c.f. [21], Theorem 3.1):
Proposition 2.3.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$ with discrete Neumann spectrum. Then for any $\lambda\geq 0$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))>L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Proof.
Again we first sketch Laptev’s proof of the corresponding (non-strict) inequality in [21]. Let $\{\psi_{k}\}_{k=0}^{\infty}$ be the $L^{2}$ normalized eigenfunctions associated to $\{\mu_{k}(\Omega)\}_{k=0}^{\infty}$ and denote $\varphi_{\lambda}(x)=(\lambda-x)_{+}$. Then
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))=(2\pi)^{-d}|\Omega|\int_{\mathbb{R}^{d}}\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))\mathrm{d}\xi.$
Using the facts
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}=1\quad\text{and}\quad\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}=|\xi|^{2}$
and the Jensen inequality (since $\varphi_{\lambda}$ is convex), Laptev proved
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))\geq\varphi_{\lambda}(|\xi|^{2})$
and thus
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))\geq L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
To show that the inequality is strict, again we fix $\lambda$ and let $N$ be the integer with $\mu_{N}(\Omega)\leq\lambda<\mu_{N+1}(\Omega)$. Take $\xi$ with $|\xi|^{2}<\lambda$ such that
$\langle e_{\xi},1\rangle_{L^{2}(\Omega)}\neq 0,\text{ and }e_{\xi}\notin\mathrm{span}\{\psi_{1},\cdots,\psi_{N}\},$
then for such a $\xi$, Jensen’s inequality is strict:
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))$
$=$ $\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))+\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))$
$=$ $(\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)})\cdot\varphi_{\lambda}(\frac{\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)}{\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}})+$
$(\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)})\cdot\varphi_{\lambda}(\frac{\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)}{\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}})$
$>$ $\varphi_{\lambda}(\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)+\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega))$
$=$ $\varphi_{\lambda}(|\xi|^{2}).$
This completes the proof. ∎
As a consequence,
Corollary 2.4.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$ with discrete Neumann spectrum. Then for any $\lambda\geq 0$ and $\gamma\geq 1$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}>L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}.$
In particular, for any $0<a<b$, one has
$K_{1}(a,b,\Omega):=\inf_{a\leq\lambda\leq b}\frac{\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}}{\lambda^{\gamma+\frac{d-1}{2}}}>0.$
@@ 2.2. Strict inequalities for Riesz means
As observed by P. Freitas, J. Lagace and J. Payette in [12, Proposition 3.1], it is enough to assume that both $\Omega_{1}$ and $\Omega_{2}$ are connected. We divide the proof of Theorem 1.1 into three parts: the Dirichlet case with $d_{2}\geq 3$, the Dirichlet case with $d_{2}=2$, and the Neumann case. For the Dirichlet case, the eigenvalues of $a\Omega_{1}\times\Omega_{2}$ are
@@
By inequality (2.8), there exists a constant $C(\Omega_{1})>0$ such that if $a^{2}\lambda>C(\Omega_{1})$, then
@@ 3.1. The Dirichlet case with $d_{2}\geq 3$
By Corollary 2.2, one has
(3.3) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}<L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
@@
one has that if $a^{2}\lambda>C(\Omega_{1})$, then
@@
then for any $\lambda>a^{-2}C(\Omega_{1})$, we will get the demanded inequality
@@
So it remains to consider the case $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$ assuming $C(\Omega_{1})>\lambda_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by Corollary 2.2, one has
$K(\Omega_{1})=\inf_{\lambda_{1}(\Omega_{1})\leq\mu\leq C(\Omega_{1})}\frac{L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}}{2}}-\sum_{\lambda_{l}(\Omega_{1})<\mu}(\mu-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}}{\mu^{\frac{d_{1}+d_{2}-1}{2}}}>0$
@@
for all $\lambda_{1}(\Omega_{1})\leq\mu\leq C(\Omega_{1})$. Thus by (3.1), (3.3) and (3.4), one has that if $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$, then
@@
So if we assume $a<\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}$, then for any $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$,
@@
$a<\min\bigg{(}\frac{C_{d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{5|\Omega_{1}|C(\Omega_{2})},\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}\bigg{)},$
then all Dirichlet eigenvalues of $a\Omega_{1}\times\Omega_{2}$ satisfy Pólya’s conjecture (2.1). This complete the proof of Theorem 1.1, part (1), for the case of $d_{2}\geq 3$.

### v3

@@ 2. Some preparations
where $|\partial\Omega|$ is the surface area of $\partial\Omega$.
Although Weyl’s conjecture was not proven in its full generality, R. Seeley ([41], [42]) proved a weaker version, namely both eigenvalue counting functions satisfy
(2.6) $\mathcal{N}^{D/N}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}+\mathrm{O}(\lambda^{\frac{d-1}{2}}),\qquad\text{as\ }\lambda\to\infty,$
@@
We also need two-term inequalities for the Riesz mean that sharpen Laptev’s inequalities (1.7) and (1.9). For the Dirichlet case, R. Frank and S. Larson ([12, Theorem 1.1]) proved that for any bounded domain $\Omega$ in $\mathbb{R}^{d}$ ($d\geq 2$) with Lipschitz boundary and any $\gamma>0$,
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}}),$
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}+\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}}),$
@@
(2.9) $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}\leq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{5}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}},$
and
@@
Remark 2.1.
If $\Omega$ is convex, R. Frank and S. Larson ([12, Theorem 1.2]) provided a uniform, non-asymptotic bound that depends on $\Omega$ only through the simple geometric characteristics. Specifically, assuming $\gamma\geq 1$ for simplicity, they proved:
$|\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}+\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}|$
$\leq$ $C(\gamma,d)|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}(r_{\mathrm{in}}(\Omega)\sqrt{\lambda})^{-\frac{1}{11}}$
and
$|\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}|$
$\leq$ $C(\gamma,d)|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}[(1+\ln_{+}(r_{\mathrm{in}}(\Omega)\sqrt{\lambda}))^{-\gamma}+(r_{\mathrm{in}}(\Omega)\sqrt{\lambda})^{1-d}]$
where $r_{\mathrm{in}}(\Omega)$ denotes the inradius of $\Omega$. By the above inequalities, $C_{1}(\Omega)$ can be chosen as a constant depending only on $r_{\mathrm{in}}(\Omega)$, $d$ and $\gamma$.
A third ingredient is a sharpened version of Laptev’s inequality (1.7) and (1.9), which is needed for us to handle eigenvalues that are neither very large nor very small. By carefully analyzing Laptev’s proof in [26], it is not hard to show that both inequalities are strict. Although this observation is enough to prove our theorem, it would be better to use an improved version so that one can say more on the constant in our theorem. In fact improvements of various forms have been obtained by many authors, see e.g. [34], [19], [20], [27], [29], [44]. What we will use below is the following quantitative improvements of both inequalities obtained recently by R. Frank and S. Larson [13], if $\gamma\geq 1$,
(2.11) $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}\leq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}(1-c\exp(-c^{\prime}\omega(\Omega)\sqrt{\lambda}))$
and
(2.12) $\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}\geq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}(1+c\exp(-c^{\prime}\omega(\Omega)\sqrt{\lambda}))$
where $c$, $c^{\prime}$ are two uniform constants and $\omega(\Omega)$ is the width of $\Omega$. For instance, by (2.11) and (2.12), one gets
$\frac{L_{\gamma,d}|\Omega|\lambda^{\frac{d}{2}+\gamma}-\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}}{\lambda^{\frac{d-1}{2}+\gamma}}\geq L_{\gamma,d}|\Omega|c\exp(-c^{\prime}\omega(\Omega)\sqrt{\lambda})\sqrt{\lambda},$
$\frac{\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}}{\lambda^{\frac{d-1}{2}+\gamma}}\geq L_{\gamma,d}|\Omega|c\exp(-c^{\prime}\omega(\Omega)\sqrt{\lambda})\sqrt{\lambda}.$
As a result, if we write $f(x)=L_{\gamma,d}|\Omega|c\exp(-c^{\prime}\omega(\Omega)\sqrt{x})\sqrt{x}$, then
(2.13) $\inf_{A\leq\lambda\leq B}\frac{L_{\gamma,d}|\Omega|\lambda^{\frac{d}{2}+\gamma}-\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}}{\lambda^{\frac{d-1}{2}+\gamma}}\geq\inf\{f(A),f(B)\},$
$\inf_{A\leq\lambda\leq B}\frac{\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}}{\lambda^{\frac{d-1}{2}+\gamma}}\geq\inf\{f(A),f(B)\}.$
@@
As observed by P. Freitas, J. Lagace and J. Payette in [15, Proposition 3.1], it is enough to assume that both $\Omega_{1}$ and $\Omega_{2}$ are connected. We divide the proof of Theorem 1.1 into three parts: the Dirichlet case with $d_{2}\geq 3$, the Dirichlet case with $d_{2}=2$, and the Neumann case.
For the Dirichlet case, the eigenvalues of $a\Omega_{1}\times\Omega_{2}$ are
@@
By inequality (2.9), there exists a constant $C_{1}(\Omega_{1})>0$ such that if $a^{2}\lambda>C_{1}(\Omega_{1})$, then
@@ 3.1. The Dirichlet case with $d_{2}\geq 3$
By (1.7), one has
(3.3) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}\leq L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
@@
one has that if $a^{2}\lambda>C_{1}(\Omega_{1})$, then
@@
then for any $\lambda>a^{-2}C_{1}(\Omega_{1})$, we will get the demanded inequality
@@
So it remains to consider the case $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C_{1}(\Omega_{1})$ assuming $C_{1}(\Omega_{1})>\lambda_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by (2.11), one has
$K(\Omega_{1})=\inf_{\lambda_{1}(\Omega_{1})\leq\mu\leq C_{1}(\Omega_{1})}\frac{L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}}{2}}-\sum_{\lambda_{l}(\Omega_{1})<\mu}(\mu-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}}{\mu^{\frac{d_{1}+d_{2}-1}{2}}}>0$
@@
for all $\lambda_{1}(\Omega_{1})\leq\mu\leq C_{1}(\Omega_{1})$. Thus by (3.1), (3.3) and (3.4), one has that if $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C_{1}(\Omega_{1})$, then
@@
So if we assume $a<\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}$, then for any $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C_{1}(\Omega_{1})$,
@@
$a<a_{0}^{D}(\Omega_{1},\Omega_{2})=\min\bigg{(}\frac{C_{d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{5|\Omega_{1}|C(\Omega_{2})},\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}\bigg{)},$
then all Dirichlet eigenvalues of $a\Omega_{1}\times\Omega_{2}$ satisfy Pólya’s conjecture (2.1).
Remark 3.1.
By (2.13), one can express $K(\Omega_{1})$ via $\lambda_{1}(\Omega_{1})$ and $C_{1}(\Omega_{1})$. In particular, if $\Omega_{1}$ is convex, then by Remark 2.1, one gets a formula of $C_{1}(\Omega_{1})$ via $r_{\mathrm{in}}(\Omega_{1})$ and $d_{1},d_{2}$. Together with the classical Faber-Krahn inequality
$\lambda_{1}(\Omega_{1})\geq\big{(}\frac{\omega_{d_{1}}}{|\Omega_{1}|}\big{)}^{\frac{2}{d_{1}}}\lambda_{1}(B^{d_{1}}),$
we may write down an explicit formula for $K(\Omega_{1})$ in terms of $|\Omega_{1}|$, $r_{\mathrm{in}}(\Omega_{1})$, $d_{1}$ and $d_{2}$ (here we used the fact that for convex domains, $\omega(\Omega_{1})$ can be controlled by $r_{\mathrm{in}}(\Omega_{1})$, see [40, Theorem 10.12.2]). As a result, the value of $a_{0}^{D}(\Omega_{1},\Omega_{2})$ can be explicitly determined if the value of $C(\Omega_{2})$ is known (See §7 for a class of such domains).